• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:04:30 on July25,2005

    Entry number 149426

    keyword=Day shift summary

    Day shift summary:
    Shift workers:
    J. P. Chen, L. Glesener , B. Reitz
    
    8:00 Right Septum went down before the shift started.
    8:00-8:30 beam off to have right Septum cool down.
    8:30 Per Run Coordinator's instruction, will take data with left arnm only 
    until scheduled down (~10 am). Asked for 35 uA.
    8:34- 10:00 data taking with left arm only.
    9:11-9:21 run 3817, adjusted left detector HV from 1650 to 1615 V to keep the 
    detector counting at 40,000.
    10:00 Beam off for injector work.
    10:10 Right Septum up coil temerature dropped below 7.8 K. Start to ramp 
    current up.
    10:20 Hall A in controlled access for target and Compton work.
    11:05 Right Septum back to 278.22A.
    16:00 beam is still off, Hall is still in controlled access. People
    are all out. 
    
    Left arm only runs:
    Run 3816, 8:34-9:11,66.6K, right septum down, 35 uA.
    Run 3817, 9:11-9:21, 16K, left HV adjustment, right septum down.
    run 3818, 9:22-10:00, 68K. Right septum down, 35 uA.
    
    DAQ test runs for Bob: 2274-2275.